a ball of thermal capacity 10cal/c is heated to the temperature of furnaace. it is then transferred into a vessel containing water . the water equivalent of vessel and the contents is 200gm . the temperature of the vessel and its contents  rises from 10c to40c . what is the temperature of furnace.

Grade:11

4 Answers

akhil
9 years ago
it is more obvious that you have problem in WATER EQUIVALENT so here it is
It is the amount of water which requires the same amount of heat for the same temperature rise as that of the object
ms T = mw Sw T
 mw =ms/sw
In calorie sw = 1
mw = ms
mw is also represent by W
so W = ms.
so if u have answer of this question i will also give u solution of this question..
Reply?
akhil
9 years ago
it is more obvious that you have problem in WATER EQUIVALENT so here it is
It is the amount of water which requires the same amount of heat for the same temperature rise as that of the object
ms T = mw Sw T
 mw =ms/sw
In calorie sw = 1
mw = ms
mw is also represent by W
so W = ms.
so if u have answer of this question i will also give u solution of this question..
Reply?
akhil
9 years ago
it is more obvious that you have problem in WATER EQUIVALENT so here it is
It is the amount of water which requires the same amount of heat for the same temperature rise as that of the object
ms T = mw Sw T
 mw =ms/sw
In calorie sw = 1
mw = ms
mw is also represent by W
so W = ms.
so if u have answer of this question i will also give u solution of this question..
Reply?
akhil
9 years ago
it is more obvious that you have problem in WATER EQUIVALENT so here it is
It is the amount of water which requires the same amount of heat for the same temperature rise as that of the object
ms T = mw Sw T
 mw =ms/sw
In calorie sw = 1
mw = ms
mw is also represent by W
so W = ms.
so if u have answer of this question i will also give u solution of this question..
Reply?